# Towards Geometric-Photometric Joint Alignment for Facial Mesh Registration

### v1

@@ 4.2 Multiscale Regularized Optimization
The HRA module guides deformation towards the joint alignment, but its derivative vectors are noisy due to shading and imaging variations. The lack of regularization of the derivatives could cause topological errors if directly applied as update steps to each vertex [30, 19]. To address these issues, a common approach is to add a smoothness term [41, 10], such as the ARAP energy [39] or the Laplacian differential representation [40]. However, these solutions introduce problems in tuning the regularization weight for outputs with both smooth and non-smooth regions [30], and implementing a robust solution scheme for non-linear optimization [38]. To overcome these challenges, we propose a multiscale regularized mechanism for generating high-quality aligned meshes, comprising two elements:
Vertex Optimization. We follow the work of Nicolet et al. [30] to update vertices iteratively. The authors suggest that the second-order optimization like Newton’s method is better for smoothing geometry, and the computationally expensive Hessian matrix can be replaced by reparameterization of $\boldsymbol{x}$ with the introduced variables $\boldsymbol{\mu}$ to ensure the smoothness of recovered $\boldsymbol{x}$:
$\boldsymbol{x}=\left(\boldsymbol{E}+\lambda\boldsymbol{L}\right)^{-1}\boldsymbol{\mu},$ (7)
@@
where $\eta>0$ means the learning rate, $\boldsymbol{E}\in\mathbb{I}^{n\times n}$ denotes identity matrix, and $\lambda>0$ is the regularization weight. $\boldsymbol{L}\in\mathbb{R}^{n\times n}$ is a discrete Laplacian operator defined on a mesh $\mathcal{M}=\left(\mathcal{X},\mathcal{E}\right)$ with $n$ vertices $\mathcal{X}$ and $m$ edges $\mathcal{E}$:
$\boldsymbol{L}_{ij}=\begin{cases}-w_{ij},&\textrm{if}\quad(i,j)\in\mathcal{E}\\ \sum_{(i,k)\in\mathcal{E}}w_{ik}&\mathrm{if}\quad i=j\\ 0&\mathrm{otherwise},\end{cases}$ (9)
@@
Combining Eq. 7 and Eq. 8, the update formula for vertices at each iteration is:
@@
$\boldsymbol{x}\leftarrow\boldsymbol{x}-\eta\left(\boldsymbol{E}+\lambda\boldsymbol{L}\right)^{-2}\frac{\partial\mathcal{L}}{\partial\boldsymbol{x}}.$ (10)
@@
Multiscale Learning. The coarse-to-fine multiscale learning scheme, based on the tessellation technique [7], periodically decreases the average edge length of the triangle mesh while retaining the shape and the texture parameterization. The multiscale scheme allows for parameter adjustment at each scale to capture fine details without distorting the topology. In particular, the Laplacian matrix $\boldsymbol{L}$ is updated for each tessellation step as the topology changes. The template mesh $\boldsymbol{T}$, initially at the coarest level, undergoes more rigid deformations with a higher regularization parameter $\lambda$ to fit the overall target expressions. As the mesh is tessellated to finer scales, $\lambda$ is decreased to capture more details.
@@
The multiscale regularized optimization offers several advantages: (1) it produces high-quality meshes effectively with significantly reduced distortion and self-intersection artifacts; (2) it converges quickly without additional training data or priors other than the textured template.
@@
Figure 4: The pipeline of the textured template mesh creation. (a) A genus-zero mesh in the shape of a bust that approximately overlaps the scan undergoes GPJA as Fig. 2, producing (b) a densely tessellated mesh accompanied with the reconstructed texture map $\boldsymbol{C}_{\boldsymbol{T}}$, which is then decimated to construct (c) a coarse template $\boldsymbol{T}$, while still preserving the same texture map.
@@
Table 1: Quantitative evaluations on 10 expressions per subject. Left Table: Our method outperforms the deep learning method NPHM [16], and the ICP-based registration FaceScape TU [43] in geometric accuracy. Right Table: The image metrics are comparable to the photo-realistic results by the NeRF-style pipeline NeP [26].
Geometric Metric(mm)↓
GPJA NPHM FaceScape TU
Subject 7 0.254 0.338 0.633
Subject 122 0.265 0.294 0.817
Subject 212 0.101 0.436 0.948
Subject 340 0.142 0.253 0.747
Subject 344 0.161 0.291 0.831
Subject 350 0.160 0.436 0.811
Overall 0.181 0.341 0.797
Image Metric
@@
24.75 0.7810 0.06868
23.56 0.7443 0.07081
25.51 0.7699 0.09368
24.10 0.7378 0.06802
23.08 0.7569 0.06791
25.38 0.7320 0.07857
24.40 0.7538 0.07461
4.3 Textured Template Mesh Creation
To adapt to the proposed pipeline, we construct a reliable textured template mesh which is used throughout the joint alignment for different expressions per subject. As illustrated in Fig. 4, we select the mouth-open expression from each subject to reconstruct it into the template mesh. This is because the mouth socket is crucial for accommodating flexible movements during deformation. Initially, we manually sculpt a coarse genus-zero mesh with pre-designed texture parametrization, resembling a bust sculpture. This genus-zero surface is appropriate for representing the head geometry due to similar topological structures without loss of generality .
The genus-zero bust mesh as initialization is processed through the GPJA pipeline depicted in Fig. 2. At this stage, only the depth constraint in HRA is involved. After the depth guiding reconstruction, a tessellated mesh is established. The vertex positions are then fixed, and the texture color map $\boldsymbol{C}_{\boldsymbol{T}}$ is updated through the color constraint via gradient descent. Finally, the tessellated mesh is decimated using seam-aware simplification [24] while preserving the texture parameterization, resulting in the creation of the textured template mesh $\boldsymbol{T}$.
5 Experiments
Experiment Setup. As an early exploration of semantic annotation-free photometric alignment, many existing public datasets (LYHM[9], NPHM[16],etc.) that primarily consist of 3D scans or registered meshes rather than original images are unsuitable for GPJA. We also found FaMoS[6] inappropriate due to its sparse down-sampled RGB views and subjects with noticeable facial markers. Following our investigation, the FaceScape dataset emerged as the most fitting benchmark with high-resolution images from dense viewpoints and uniform lighting conditions for the subjects.
In order to thoroughly assess GPJA’s capability, six subjects including four publishable ones are chosen, and we deliberately selected 10 highly different expressions with large deformation and occlusion variations for each subject to evaluate GPJA. Six to eight images covering frontal and side views are used as color reference images, and are downsampled to 2K resolution.
GPJA is implemented using PyTorch with Nvdiffrast [20] as the differentiable renderer. The multiscale scheme involves remeshing for 4 times, increasing the vertex count from 16K to 250K. For the first two levels, the regularization parameter $\lambda$ is set to 200 and 120, while for the remaining levels, it is set to 80 and 50. Convergence is achieved within 1500 iterations per expression, taking approximately 15 minutes on a single RTX 3090 graphics card for all experiments.
Metrics. We evaluate the effectiveness of the method from both geometric and photometric aspects, since GPJA achieves joint alignments. The assessment of geometric alignment uses raw scans of the face region as ground truth, measuring distances between ground truth vertices and their closest Euclidean distance on aligned surfaces in millimeter. For photometric alignment assessment, multi-view images captured by the setup are considered as ground truth, while rendered images are synthesized at the same camera poses using aligned meshes with the texture color map of the template. We utilize commonly used image metrics including PSNR, SSIM, and LPIPS for evaluation.
Figure 5: Visualization on the geometric errors in point-to-mesh distance. The geometric error of our method is significantly lower than NPHM and FaceScape TU meshes.
(a) Inconsistent textures of FaceScape TU
(b) False lips by NPHM
Figure 6: Qualitative evaluation on the photometric alignment of FaceScape TU and NPHM meshes. Both registration methods only take into account the geometry, resulting in obvious photometric misalignment.
@@ 5 Experiments
Result Analysis on Geometric Alignment. We compare our GPJA against two registration methods. The first is FaceScape topologically uniform(TU) meshes obtained through a variant method of ICP [43] which is the standard surface registration method in 3D face domain [19]. The second is the-state-of-the-art deep learning based method NPHM [16], which is trained on a dataset comprising 87 subjects and 23 different facial expressions.
Quantitative and qualitative comparisons of geometric alignment are presented in the left of Tab. 1 and Fig. 5. The statistical analysis demonstrates the overall superior performance of our method. As illustrated in Fig. 5, our method achieves higher fidelity, even in challenging areas such as the lips and eyes, and captures finer details of wrinkles on the jaws and cheeks. The primary limitation of NPHM lies in the lack of fidelity in capturing details. This deficiency is particularly evident in errors concentrated around the mouth of subjetc 212 and wrinkles of subject 344.  Figure 5 also identifies two main drawbacks in FaceScape TU meshes. Firstly, subject 122’s lip-funneler expression, where the eyes should be closed, exhibits the opposite. This is a common issue resulting from inaccurate landmarks in the ICP-based method. Secondly, the face rim areas (forehead and cheeks) in TU meshes are observed with higher geometric errors due to excessive conformation to the template shape, which is related to the issue of selecting suitable parameters for appropriate global settings.
Table 2: The ablation studies to demonstrate the effects of each constraint from HRA. The geometric performance degrades when one of three constraints is not equipped, validating the contributions of the three constrains in our GPJA.
Geometric Metric↓ PSNR↑ SSIM↑ LPIPS↓
@@
(a) Ablations with and without
the normal constraint
(b) Ablations with and without
the mask operation
Figure 8: Ablation studies. (a) The normal constraint in HRA significantly enhanced details. (b) Masking out the interior mouth avoids disturbance around the contour.
Result Analysis on Photometric Alignment. Previous registration methods are prone to neglect photometric information, resulting in a failure to consistently preserve texture parameterization. The FaceScape TU meshes are accompanied by texture maps for each expression. Nevertheless, inconsistencies in the texture maps of the same subject across various expressions are commonly observed (Fig. 5(a)), illustrating the failure of TU meshes in maintaining photometric alignment. Deep learning methods trained on aligned meshes processed by ICP-based techniques also struggle to address this issue. Notably, Fig. 5(b) reveals another common drawback due to the overlook of photometric features: in particular, NPHM outputs false lips that should be invisible. While this issue is obvious in photometric appearances, it is not apparent in geometric evaluations.
On the contrary, in GPJA, photometric consistency is ensured through applying the shared color texture to fit reference images of diverse expressions. Comparison between the ground truth and our rendered images in Fig. 7 convincingly demonstrates the photometric alignment achieved through our registration. Notably, the rendered images exhibit pixel-level overlapping consistency in characteristic areas such as the mouth, eyes, and mole features, across various facial expressions despite occlusion changes. The last row of Fig. 7 illustrates that the discrepancy between the ground truth and the rendered images is primarily attributed to variations in skin tone across facial expressions. Additionally,the right of Tab. 1 provides the image metrics for our experiment. The quantitative results are superior than PSNR of 23.61, SSIM of 0.6460, and LPIPS of 0.09677 achieved by the NeRF-style pipeline NeP [26] (tested on the first 100 subjects of FaceScape), which produces photo-realistic reconstruction results through per-frame color generation.
@@
Ablations. The HRA based on multiple cues plays a crucial role in correctly warping the template into joint alignment. To validate HRA’s effectiveness, we conduct two ablation experiments to confirm its benefits utilizing four publishable subjects of FaceScape.
@@
We first validate the contribution of each constraints from HRA mechanism, which is supported by  Tab. 2. In the case where all constraints are utilized, the geometric error is minimized. As a visualized example, Fig. 7(a) reveals that the normal term significantly contributes to the sharpness of details and alleviates incorrect bumps on texture-less regions like cheeks. However, Tab. 2 show close image metrics in comparison. We speculate the reason is that with the color constraint guiding the deformation, the geometric distortion can not manifest itself in color renderings.
@@
In the second ablation, we study the effectiveness of the masking mechanism for color constraint. We synthesize the label image for the mouth socket using $F_{S}(\boldsymbol{x}|\boldsymbol{P}_{j},\boldsymbol{B})$, and overlap it with the ground truth to examine the contours around the inner mouth. Figure 7(b) illustrates that when the inner mouth is not masked out, the vertices around it are disturbed and become distorted. Hence, intentionally masking out the inner mouth facilitates correct tracing of the mouth contours.

### v2

@@ 4.2 Multiscale Regularized Optimization
The HRA mechanism steers deformation towards joint alignment, yet its derivative vectors are noisy because of shading and imaging variations. Since the derivatives lack regularization, applying them directly as update steps to each vertex could lead to topological errors [17, 57]. A common way to tackle this is adding a regularization term [28, 56], like the ARAP energy [58] or the Laplacian differential representation [19]. However, these solutions introduce problems in tuning the regularization weight for outputs with both smooth and non-smooth regions [17], and implementing a robust solution scheme for non-linear optimization [59]. To overcome these challenges, we propose a multiscale regularized optimization for generating high-quality aligned meshes.
Vertex Optimization. We follow the work of [17] to update vertices iteratively. It suggests that the second-order optimization like Newton’s method is better for smoothing geometry, and the computationally expensive Hessian matrix can be replaced by re-parametrization of $\boldsymbol{x}$ with the introduced variables $\boldsymbol{\mu}$ to ensure the smoothness of recovered $\boldsymbol{x}$:
$\boldsymbol{x}=\left(\boldsymbol{I}+\lambda\boldsymbol{L}\right)^{-1}\boldsymbol{\mu},$ (7)
@@
where $\eta>0$ means the learning rate, $\boldsymbol{I}\in\mathbb{I}^{n\times n}$ denotes identity matrix, and $\lambda>0$ is the regularization weight. $\boldsymbol{L}\in\mathbb{R}^{n\times n}$ is a discrete Laplacian operator defined on a mesh $\mathcal{M}=\left(\mathcal{X},\mathcal{E}\right)$ with $n$ vertices $\mathcal{X}$ and $m$ edges $\mathcal{E}$:
$\boldsymbol{L}_{ij}=\left\{\begin{matrix}-w_{ij},&\textrm{if}\quad(i,j)\in\mathcal{E}\\ \sum_{(i,k)\in\mathcal{E}}w_{ik}&\mathrm{if}\quad i=j\\ 0,&\mathrm{otherwise},\end{matrix}\right.\$ (9)
@@
Combining Equations 7 and 8, the updated formula for vertices at each iteration is:
@@
$\boldsymbol{x}\leftarrow\boldsymbol{x}-\eta\left(\boldsymbol{I}+\lambda\boldsymbol{L}\right)^{-2}\frac{\partial\mathcal{L}_{HRA}}{\partial\boldsymbol{x}}.$ (10)
@@
Multiscale Learning. The coarse-to-fine multiscale learning scheme, based on the tessellation technique [61], periodically decreases the average edge length of the triangle mesh while retaining the shape and the texture parametrization. The multiscale scheme allows for parameter adjustment at each scale to capture fine details without distorting the topology. In particular, the Laplacian matrix $\boldsymbol{L}$ is updated for each tessellation step as the topology changes. The template mesh $\boldsymbol{T}$, initially at the coarsest level, undergoes more rigid deformations with a higher regularization parameter $\lambda$ to fit the overall target expressions. As the mesh is tessellated to finer scales, $\lambda$ is decreased to capture more details.
@@
The multiscale regularized optimization offers several advantages: (1) it produces high-quality meshes effectively with significantly reduced distortion and self-intersection artifacts; (2) it converges robustly without additional training data or priors other than the textured template. The iterative optimization process can be observed in the supplementary videos.
@@
Figure 5: The pipeline of the textured template mesh creation. (a) A genus-0 mesh in the shape of a bust that approximately overlaps the scan undergoes GPJA as Fig. 2, producing (b) a densely tessellated mesh accompanied with the reconstructed texture map $\boldsymbol{C}_{\boldsymbol{T}}$, which is then decimated to construct (c) a coarse template $\boldsymbol{T}$, while still preserving the same texture map.
4.3 Textured Template Mesh Creation
To adapt to the proposed pipeline, we construct a reliable textured template mesh which is used throughout the joint alignment for different expressions per subject. As illustrated in Fig. 5, we deliberately choose the mouth-open expression from each subject to reconstruct it into the template mesh. This is because the mouth socket is crucial for accommodating flexible movements during deformation. Initially, we manually sculpt a coarse genus-0 mesh with pre-designed texture parametrization, resembling a bust sculpture. This genus-0 surface is appropriate for representing the head geometry due to similar topological structures without loss of generality .
The genus-0 bust mesh as initialization is then processed through the GPJA pipeline. At this stage, only the depth constraint in HRA is applied. After the depth guiding reconstruction, a tessellated mesh is established. The geometry is then fixed, and the texture color map $\boldsymbol{C}_{\boldsymbol{T}}$ is updated through the color constraint via gradient descent, shown in Fig. 5(b). Finally, the tessellated mesh is decimated using seam-aware simplification [62] while preserving the texture parametrization, resulting in the creation of the textured template mesh $\boldsymbol{T}$ as the startup mesh in Fig. 2.
Although the template is subject-specific, all expressions share a common texture space, which enables batch re-parametrization. This offers a practical way for topology-consistent meshes of large datasets.
5 Experiments and Analysis
Experiment Setup. As an early exploration of semantic annotation-free photometric alignment, many existing public datasets (LYHM [25], and NPHM [41], etc.) that primarily consist of 3D scans or registered meshes rather than original images are unsuitable for GPJA. We also found FaMoS [40] inappropriate due to its sparse down-sampled RGB views and subjects with noticeable facial markers. Following our investigation, the FaceScape dataset emerged as the most fitting benchmark with high-resolution images from dense viewpoints and uniform lighting conditions for discrete facial expressions.
In order to thoroughly assess GPJA’s capability, seven subjects, including four publishable ones (Subject 122, 212, 340 and 344), are chosen, and we deliberately selected 10 highly different expressions for each subject. The selection process eliminates redundant and similar expressions. Additionally, we prioritize expressions with significant deformation and occlusion compared to the neutral expression, where landmark detection is less accurate as shown in Fig. 3. This process ensures that the chosen expressions cover a range of challenging scenarios and effectively evaluate the algorithm’s performance. Six to eight images, covering frontal and side views, are used as references.
GPJA is implemented with Nvdiffrast [12] as the differentiable renderer and LargeSteps [17] as the optimizer. The rendering uses the Lambertian reflection model under uniform lighting, which is a reasonable simplification of FaceScape’s acquisition environment. The multiscale optimization involves remeshing 4 times, increasing the vertex count from 16K to 250K. The learning rate $\eta$ is set to one-tenth of the face length. For the first two levels, the regularization parameter $\lambda$ is set to 200 and 120, while for the remaining levels, it is set to 80 and 50. Gradients of only one constraint are back-propagated per iteration, with three constraints being iterated in rotation. Convergence is achieved within 1500 iterations per expression, taking approximately 15 minutes on a single NVIDIA RTX 3080 graphics card.
Metrics. As GPJA achieves joint alignment, we evaluate the method’s effectiveness from both geometric and photometric perspectives. Geometric alignment is assessed using raw face scans as ground truth, with L1-Chamfer distance and normal consistency reported. The distance is computed via the point-mesh approach to reduce the influence of vertex counts. Moreover, we calculate the F-Score with thresholds of 0.5mm and 1.0mm for a statistical error analysis. For photometric alignment, multiview reference images are used as ground truth, while rendered images are generated from aligned meshes with the common template’s texture map at the same camera poses. Standard image metrics such as PSNR, SSIM, and LPIPS are employed for evaluation.
Table 1: Geometric evaluations on 10 expressions per subject in millimeter. C.D. stands for L1-Chamfer distance, N.C. represents normal consistency, while F@0.5 and F@1.0 are F-Scores at thresholds of 0.5mm and 1.0mm.
Subject 7 Subject 32 Subject 122 Subject 212 Subject 340 Subject 344 Subject 350
C.D.$\downarrow$ N.C.$\uparrow$ C.D.$\downarrow$ N.C.$\uparrow$ C.D.$\downarrow$ N.C.$\uparrow$ C.D.$\downarrow$ N.C.$\uparrow$ C.D.$\downarrow$ N.C.$\uparrow$ C.D.$\downarrow$ N.C.$\uparrow$ C.D.$\downarrow$ N.C.$\uparrow$
NICP 0.477 0.910 0.355 0.942 0.404 0.943 0.614 0.915 0.383 0.944 0.340 0.936 0.428 0.912
NPHM 0.347 0.958 0.377 0.970 0.325 0.968 0.429 0.949 0.285 0.973 0.338 0.957 0.489 0.937
GPJA 0.229 0.981 0.147 0.992 0.264 0.984 0.180 0.980 0.215 0.981 0.224 0.975 0.239 0.970
F@0.5$\uparrow$ F@1.0$\uparrow$ F@0.5$\uparrow$ F@1.0$\uparrow$ F@0.5$\uparrow$ F@1.0$\uparrow$ F@0.5$\uparrow$ F@1.0$\uparrow$ F@0.5$\uparrow$ F@1.0$\uparrow$ F@0.5$\uparrow$ F@1.0$\uparrow$ F@0.5$\uparrow$ F@1.0$\uparrow$
NICP 0.708 0.852 0.800 0.918 0.738 0.901 0.549 0.799 0.743 0.913 0.792 0.922 0.746 0.879
NPHM 0.814 0.923 0.778 0.925 0.825 0.945 0.731 0.899 0.858 0.956 0.814 0.938 0.682 0.872
GPJA 0.909 0.960 0.957 0.975 0.880 0.959 0.921 0.963 0.912 0.961 0.912 0.960 0.891 0.957
Figure 6: Visualization on the geometric errors. GPJA outputs meshes with higher vertex counts than NPHM and FaceScape TU. To mitigate the effect of vertex counts, we use the point(of GT)-to-mesh distance. The error distribution show that GPJA achieves lower geometric errors compared to NPHM and FaceScape TU.
Result Analysis on Geometric Alignment. We compare our GPJA against two registration methods. The first is FaceScape topologically uniform(TU) meshes obtained through NICP [18] which is a variant of the standard ICP registration method in 3D face domain [57]. The second is the state-of-the-art deep learning method NPHM [41], which is trained on a dataset comprising 87 subjects and 23 different facial expressions.
Quantitative and qualitative comparisons of geometric alignment are presented in Table 1 and Fig. 6. Table 1 shows that GPJA outperforms NPHM and NICP across all subjects, achieving the lowest L1-Chamfer distance and highest normal consistency, indicating superior geometric alignment. GPJA also consistently achieves higher F-Scores at both thresholds, with F-Scores@0.5mm particularly emphasizing its statistical superiority.
As depicted in Fig. 6, the error visualization exhibits the distribution of errors across the face, demonstrating that our method achieves higher fidelity, even in challenging regions such as the lips and eyes. The primary limitation of NPHM is its lack of fidelity in capturing details. This deficiency is particularly evident in errors concentrated around the mouth of subject 212 and wrinkles of subject 344. Fig. 6 also reveals two main drawbacks in FaceScape TU meshes. First, subject 122’s lip-funneler expression, where the eyes should be closed, incorrectly shows the opposite. This is a common issue resulting from inaccurate landmarks in the ICP-based method. Secondly, the face rim areas (foreheads and cheeks) in TU meshes are observed with higher geometric errors due to excessive conformation to the template shape. Additionally, both NPHM and TU meshes incorrectly stitch non-facial areas like the skull, ears, and neck using templates. In contrast, GPJA aligns these regions more accurately, showcasing its robustness in capturing full heads.
Result Analysis on Photometric Alignment. Previous registration methods neglect photometric evaluations due to a lack of consideration for joint alignment, hence we provide qualitative evaluation for comparisons.
Figure 7: Evidences of photometric misalignment on FaceScape TU meshes.
Figure 8: Evidences of photometric misalignment in NPHM. The lip contours on the NPHM mesh do not align with the GT images.
There are two key evidences illustrating the lack of photometric alignment in FaceScape, as shown in Fig. 7: (a) The texture maps for the same subject exhibit discrepancies across different expressions, signifying the absence of a unified parametrization; (b) Deviations are observed between the contours of lips on the texture maps and those on the meshes, revealing misalignment between the geometry and texture. Similar issues are also present in NPHM meshes, although they are not accompanied by texture coordinates. Fig. 8 reveals that NPHM generates false lips that ought to be invisible, which implies the method’s insufficient ability to align based solely on geometric features.
@@ 5 Experiments and Analysis
Table 2: Photometric evaluations on 10 expressions per subject. The image metrics are comparable to the photo-realistic results by the NeRF-style reconstruction NeP [63].
@@
Subject 7 24.75 0.7810 0.06868
Subject 32 20.79 0.6621 0.07516
Subject 122 23.56 0.7443 0.07081
Subject 212 25.51 0.7699 0.09368
Subject 340 24.10 0.7378 0.06802
Subject 344 23.08 0.7569 0.06791
Subject 350 25.38 0.7320 0.07857
Average 23.88 0.7406 0.07469
@@
In contrast, GPJA ensures photometric consistency by applying the shared color texture to match reference images of different expressions. A comparison between the ground truth and our rendered images in Fig. 9 convincingly shows the photometric alignment achieved by our registration method. The last row of Fig. 9 illustrates that the discrepancy between the ground truth and the rendered images is primarily attributed to variations in skin tone across facial expressions. Notably, as depicted in Fig. 10, the rendered images exhibit pixel-level consistency in characteristic areas such as the mouth, eyes, and mole features across various facial expressions, despite occlusion changes. Moreover, Table 2 provides the image metrics for our experiment. The quantitative results are superior than PSNR of 23.61, SSIM of 0.6460, and LPIPS of 0.09677 achieved by the NeRF-style pipeline NeP [63] (tested on the first 100 subjects of FaceScape), which produces photo-realistic reconstruction through per-frame color generation without alignment.
Figure 10: Close-up views of GT and rendered images. Checkerboard patterns are used to demonstrate the photometric consistency of moles and freckles, as indicated by arrows. Zooming in on this figure is recommended.
Figure 11: Ablations with and without the normal constraint. The normal constraint in HRA significantly enhanced details.
Figure 12: Ablations with and without the mask operation. Masking out the interior mouth avoids disturbance around the contour.
Table 3: Ablation studies demonstrating the effects of each constraint from HRA on four publishable subjects. The geometric error used is the point-to-mesh distance. The geometric performance degrades when any of the three constraints is removed, validating the contributions of each constraint in HRA.
Geometric error↓ PSNR↑ SSIM↑ LPIPS↓
@@
Ablations. The HRA mechanism plays a crucial role in correctly warping the template into joint alignment. To validate HRA’s effectiveness, we conduct two ablation experiments to confirm its benefits.
@@
We first validate the contribution of each constraints from HRA mechanism, which is supported by Table 3. In the case where all constraints are utilized, the geometric error is minimized. As a visualized example, Fig. 11 reveals that the normal term significantly contributes to the sharpness of details and alleviates incorrect bumps on texture-less regions like cheeks. However, Table 3 show close image metrics in comparison. We speculate the reason is that with the color constraint guiding the deformation, the geometric distortion can not manifest itself in color renderings.
@@
In the second ablation, we study the effectiveness of the masking strategy for color constraint. We synthesize the label image for the mouth socket using $F_{S}(\boldsymbol{x}|\boldsymbol{\Pi}_{j},\boldsymbol{B})$, and overlap it with the ground truth to examine the contours around the inner mouth. Fig. 12 illustrates that when the inner mouth is not masked out, the vertices around it are disturbed and distorted. Hence, intentionally masking out the inner mouth facilitates correct tracing of the mouth contours.
Applications. The core strength of GPJA lies in its ability to achieve joint alignment with a shared texture space across expressions. This consistency enables efficient batch processing for a variety of applications, such as animation and re-parametrization, as demonstrated below.
The meshes generated by GPJA for the same subject share a common parametrization, though the vertex sampling varies. We perform remeshing on these meshes via a common triangulation on the texture space. The remeshed results can then be linearly interpolated for face animation as presented in Fig. 13 and the supplementary video.
When an artist-edited, UV-unwrapped facial mesh is available, as illustrated in Fig. 14(a), the texture embeddings of input meshes yield a continuous map $\varphi_{b}\varphi_{a}^{-1}$ from the texture coordinates of the GPJA-generated meshes to those of the artist-edited mesh. Thereafter, the interpolation operator is applied directly to propagate the artist-edited parametrization across all expressions shown in Fig. 14(b).
Figure 13: Face animation via vertex interpolation.
Figure 14: (a) Computation of maps on the texture spaces. (b) Re-parametrization results.